I need to get 500 gallons of a 34% mixture but I only have 5oo gallons of a 25% mixture and 500 gallons of a 50% mixture what is the formula to mix? Thank you

Answer by stanbon(75887)

I need to get 500 gallons of a 34% mixture but I only have 5oo gallons of a 25% mixture and 500 gallons of a 50% mixture what is the formula to mix? Thank you
---------------------------------
Equation:
Quantity Equation:::::x + y = 500
Active Ingredient Eq::0.25x + 0.50y = 0.34*500
---------------------------------
25x + 25y = 25*500
25x + 50y = 34*500
------------------------------
25y = 9*500
y = 180 gallons (amt of 50% solution needed in the mixture)
----
x + y = 500
x + 180 = 500
x = 320 gallons (amt of 25% solution needed in the mixture)
===============================
Cheers,
Stan H.